# RoVerFly: Robust and Versatile Learning-based Control of Quadrotor Across Payload Configurations

### v1

@@ IV-C1 State
Position and velocity vectors are normalized element-wise by their respective maxima $x_{Q}^{\max}$, $x_{P}^{\max}$, $v_{Q}^{\max}$, and $v_{P}^{\max}$. Attitude is represented by a rotation matrix to avoid singularities and remain compatible with geometric-control formulations.
IV-C2 Action
@@
A smooth squashing (e.g., $\tanh$) may be used; in practice we use linear rate scaling and an affine map for $f_{t}$. The desired rates $\boldsymbol{\Omega}^{\,d}_{t}$ are tracked by a per-axis PPID acting on $\boldsymbol{\Omega}_{t}$ with anti-windup and saturation, yielding $\mathbf{M}_{t}\in\mathbb{R}^{3}$; this inner loop decouples attitude from thrust and is robust to delays and model mismatch. Finally, $(f_{t},\mathbf{M}_{t})$ is mapped to single-rotor thrusts via a fixed mixer $\mathbf{A}\in\mathbb{R}^{4\times 4}$ set by geometry and rotor coefficients,
@@
and, in simulation, passed through first-order motor dynamics. To reduce sim-to-real gap, we also inject a 10–30 ms discrete input delay at 100 Hz and apply a first-order rotor lag with asymmetric rise/fall time constants, randomized during training.
IV-C3 Observation
@@
where $\mathbf{s}_{t}$ is the current state and $\mathbf{a}_{t-1}$ the previous action.
$\mathbf{e}_{t}\in\mathbb{R}^{12}$ is the body-frame tracking error:
@@
where $\mathbf{e}_{x_{P}},\mathbf{e}_{v_{P}},\mathbf{e}_{x_{Q}},\mathbf{e}_{v_{Q}}$ are the position/velocity errors for payload and quadrotor.
$\mathbf{\zeta}_{t}$ is the I/O history:
@@
$\mathbf{\xi}_{t}$ is the body-frame feedforward term w.r.t. the previewed reference:
@@
To emulate sensor imperfections and promote robustness under partial observability, we inject clipped Gaussian noise during training: positions $\mathbf{x}_{Q},\mathbf{x}_{P}$ use $\sigma_{x}{=}0.01$ m with $\pm 2.5$ mm clipping; linear velocities $\mathbf{v}_{Q},\mathbf{v}_{P}$ use $\sigma_{v}{=}0.02$ m/s with $\pm 5$ mm/s clipping; attitude uses $\tilde{\mathbf{R}}=\mathbf{R}(\boldsymbol{\theta}+\eta_{\theta})$ with $\sigma_{\theta}{=}\pi/60$ rad and $\pm\pi/120$ rad clipping; body rates $\boldsymbol{\Omega}$ use $\sigma_{\Omega}{=}\pi/30$ rad/s with $\pm\pi/60$ rad/s clipping.
In principle, a full history would be ideal but is intractable; a fixed-length buffer offers a practical POMDP approximation that preserves essential temporal information [28]. While I/O history is common in legged locomotion [29, 30], high control rates in aerial robots (often >50>50 Hz) make compact encodings important. With a payload, the hybrid and more underactuated dynamics further increase the value of history: it aids implicit dynamics learning (analogous to model-based RL that first learns dynamics, then optimizes control [31]) and provides temporal feedback that improves robustness to delays and sensor noise [29]. Finally, organizing the observation into present (feedback/state/action), past (I/O history), and future (feedforward preview) mirrors finite-horizon optimal-control structure (e.g., MPC) and helps the policy exploit known future references when maximizing return [32].
IV-D Rewards and Episode Design
IV-D1 Reward structure
@@ IV-D2 Early termination
IV-E Control Perspective and Stability Considerations
At a high level, the learned policy issues CTBR commands while a faster rate–PPID and static mixer produce body moments and single–rotor thrusts; with anti–windup and saturation, the inner loop behaves nearly as an identity from $\boldsymbol{\Omega}^{\,d}$ to body rates.
@@ IV-E Control Perspective and Stability Considerations
(b) Normalized settling time $T_{s}/T_{n}$.
@@ IV-F Domain Randomization
IV-G Belief-Based Analysis of I/O History in Partially Observable Hybrid Systems
In partially observable systems such as the quadrotor–payload system, inferring the full state $\mathbf{s}_{t}$ from a single observation $\mathbf{o}_{t}$ is intractable due to hybrid dynamics from the flexible cable. We therefore use a fixed-length history of observations and actions (I/O history) to approximate the belief state and briefly justify this choice.
IV-G1 Belief-Based Policy and Value under Uncertainty
This update reduces entropy by exploiting temporal context:
Lower-entropy beliefs improve performance under the value function:
Hence, belief accuracy is critical in POMDPs.
Figure 3: Payload tracking on unseen trajectories: without payload (top) and with flexible cable-suspended payload (bottom).
IV-G2 Hybrid Dynamics and Mode Estimation
Flexible cables induce hybrid dynamics:
where the mode $m_{t}$ is unobserved. The agent maintains a joint belief over mode and state:
Updating the mode belief marginalizes past modes and states:
$\qquad\times b_{t-1}(m_{t-1},\mathbf{s}_{t-1})\,d\mathbf{s}_{t-1}.$
Because mode transitions depend on latent variables, a single observation $\mathbf{o}_{t}$ is insufficient; temporal context is needed to disambiguate modes. Entropy shrinks with longer histories:
$>H(m_{t},\mathbf{s}_{t}\mid\mathbf{o}_{0:t},\mathbf{a}_{0:t-1}).$
Thus, I/O history is essential for reducing uncertainty over both mode and state, and is particularly beneficial for robust control in quadrotor–payload systems with hybrid, stochastic dynamics—as corroborated by the ablation results in Table IV.
V Experiments
@@
Mass $0.280~\text{kg}$
Inertia around $x,y$ $2.36\times 10^{-4}~\text{kg}\cdot\text{m}^{2}$
Inertia around $z$ $3.03\times 10^{-4}~\text{kg}\cdot\text{m}^{2}$
Arm Length $0.058~\text{m}$
Propeller Thrust Factor $1.145\times 10^{-7}~\text{N}\cdot\text{s}^{2}$
Payload Mass $0.025~\text{kg}$
TABLE III: Quadrotor parameters.
@@ V-F2 Cable direction and quadrotor reference
Figure 5: Example payload trajectory with nominal cable-tension directions (red).
@@ VI Conclusion and Future Work
We presented RoVerFly, a unified learning-based controller that tracks arbitrary trajectories and stabilizes from disturbances across payload conditions, from no payload to flexible cable–suspended loads with varying mass and cable length. A single policy, trained with task and domain randomization, handles changes in $(m_{P},l)$ without controller switching or re-tuning. The observation design combines present feedback, short I/O history, and finite-horizon feedforward terms, and the CTBR action parameterization yields smooth attitude shaping and effective payload damping.
Experiments show (i) low tracking error across configurations, (ii) rapid disturbance rejection with near-zero steady-state error and settling times on the order of one pendulum period, (iii) smooth performance trends over $(m_{P},l)$ without brittle failures, and (iv) ablations confirming the value of history and feedforward information. Together, these results indicate strong zero-shot generalization of a single policy across hybrid, underactuated regimes.
Limitations. We assume a taut, massless cable; slack–taut transitions and contact events are only partially captured. Evaluation is in high-fidelity simulation; real-world sensing artifacts, unmodeled aerodynamics, and onboard compute limits are not yet systematically assessed. Finally, the deployed policy has no formal safety or constraint guarantees.
Future work. We will (i) validate on hardware across platforms and sensors, including vision-in-the-loop estimation; (ii) extend coverage to slack cables, contact, and time-varying cable length, and benchmark under aggressive, dynamically infeasible references; (iii) integrate safety wrappers (e.g., control-barrier-function shields) and online identification of $(m_{P},l)$ for constraint-aware adaptation; (iv) explore meta-RL and few-shot test-time adaptation for out-of-distribution payloads and environments; (v) scale to multi-UAV cooperative transport and manipulation; and (vi) develop analytical links (e.g., input-to-state stability, robustness margins) between the learned policy and classical control guarantees.

### v2

@@ IV-D1 State
Position and velocity vectors are normalized element-wise by their respective maxima $x_{Q}^{\max}$, $x_{P}^{\max}$, $v_{Q}^{\max}$, and $v_{P}^{\max}$. Attitude is represented by a rotation matrix to avoid singularities and remains compatible with geometric-control formulations. Unlike existing work that treats payload configuration as unknown parameters and only ensures policy performance within a prior parameter set, we explicitly include payload information by augmenting mass and cable length as states.
IV-D2 Action
@@
A smooth squashing (e.g., $\tanh$) may be used; in practice we use linear rate scaling and an affine map for $f_{t}$. The desired rates $\boldsymbol{\Omega}^{\,d}_{t}$ are tracked by a per-axis PID acting on $\boldsymbol{\Omega}_{t}$ with anti-windup and saturation, yielding $\mathbf{M}_{t}\in\mathbb{R}^{3}$; this inner loop decouples attitude from thrust and is robust to delays and model mismatch. Finally, $(f_{t},\mathbf{M}_{t})$ is mapped to single-rotor thrusts via a fixed mixer $\mathbf{A}\in\mathbb{R}^{4\times 4}$ set by geometry and rotor coefficients,
@@
and passed through first-order motor dynamics. To reduce sim-to-real gap, we also inject a 10–30 ms discrete input delay at 100 Hz and apply a first-order rotor lag with asymmetric rise/fall time constants, randomized during training.
IV-D3 Observation
@@
where $\mathbf{s}_{t}$ is the current state and $\mathbf{a}_{t-1}$ the previous action. $\mathbf{e}_{t}\in\mathbb{R}^{12}$ is the body-frame tracking error:
@@
where $\mathbf{e}_{x_{P}},\mathbf{e}_{v_{P}},\mathbf{e}_{x_{Q}},\mathbf{e}_{v_{Q}}$ are the position/velocity errors for payload and quadrotor. $\mathbf{\zeta}_{t}$ is the I/O history:
@@
$\mathbf{\xi}_{t}$ is the body-frame feedforward term w.r.t. the preview:
@@
To emulate sensor imperfections and promote robustness under partial observability, we inject clipped Gaussian noise during training: positions $\mathbf{x}_{Q},\mathbf{x}_{P}$ use $\sigma_{x}{=}0.01$ m with $\pm 2.5$ mm clipping; linear velocities $\mathbf{v}_{Q},\mathbf{v}_{P}$ use $\sigma_{v}{=}0.02$ m/s with $\pm 5$ mm/s clipping; attitude uses $\tilde{\mathbf{R}}=\mathbf{R}(\boldsymbol{\theta}+\eta_{\theta})$ with $\sigma_{\theta}{=}\pi/60$ rad and $\pm\pi/120$ rad clipping; body rates $\boldsymbol{\Omega}$ use $\sigma_{\Omega}{=}\pi/30$ rad/s with $\pm\pi/60$ rad/s clipping. In principle, a full history would be ideal but is intractable; a fixed-length buffer offers a practical POMDP approximation that preserves essential temporal information [27]. While I/O history is common in legged locomotion [28, 29], high control rates in aerial robots (often >50>50 Hz) make compact encodings important. With a payload, the hybrid and more underactuated dynamics further increase the value of history: it aids implicit dynamics learning (analogous to model-based RL that first learns dynamics, then optimizes control) and provides temporal feedback that improves robustness to delays and sensor noise [28]. Finally, organizing the observation into present (feedback/state/action), past (I/O history), and future (feedforward preview) mirrors finite-horizon optimal-control structure and helps the policy exploit known future references when maximizing return [30].
IV-E Rewards and Episode Design
IV-E1 Reward structure
@@ IV-E2 Early termination
IV-F Control Perspective and Stability Considerations
At a high level, the learned policy issues CTBR commands while a faster rate–PID and static mixer produce body moments and single–rotor thrusts; with anti–windup and saturation, the inner loop behaves nearly as an identity from $\boldsymbol{\Omega}^{\,d}$ to body rates.
@@ IV-F Control Perspective and Stability Considerations
(b) Normalized settling time.
@@ IV-G Domain Randomization
V Numerical Results
@@
Figure 3: Payload tracking on unseen trajectories: without payload and with flexible cable-suspended payload.
@@
$m_{Q}$ $0.835~\text{kg}$
$m_{P}$ $0.0-0.2~\text{kg}$
$J_{Q}$ $Diag([4.01,3.58,6.36])\times 10^{-3}~\text{kg}\cdot\text{m}^{2}$
$\bar{f}$ $30~\text{N}$
TABLE III: Nominal quadrotor parameters.
@@ V-F2 Cable direction and quadrotor reference
(a) Quadrotor thrust in $z$ axis.
(b) Payload trajectory in $z$.
Figure 5: Controller behavior during slack-to-taut transition. (a) The controller modulates thrust before tautening, mitigating impact. (b) The payload exhibits distinct free-fall and bounce slack phases before settling. The shaded intervals highlight free-fall and bounce phases during slack-to-taut transition.
@@ VI Conclusion and Future Work
We presented RoVerFly, a unified learning-based controller that tracks arbitrary trajectories and stabilizes under disturbances across payload conditions, from no payload to flexible cable–suspended loads with varying mass and length. A single policy trained with task and domain randomization handles $(m_{P},l)$ changes without switching or re-tuning. The observation design combines feedback, short I/O history, and feedforward terms, while CTBR parameterization yields smooth attitude shaping and effective payload damping. Experiments show (i) low tracking error across configurations, (ii) rapid disturbance rejection with near-zero steady-state error and settling near one pendulum period, (iii) smooth trends over $(m_{P},l)$ without brittle failures, and (iv) ablations confirming the role of history and feedforward, indicating strong zero-shot generalization of a single policy across hybrid, underactuated regimes.
Our analysis in uses the standard taut, massless cable model, while simulations employ a flexible cable with slack–taut transitions. The learned policy handles these empirically, but stability guarantees for the full hybrid system remain open. Evaluation is limited to high-fidelity simulation; unmodeled aerodynamics, and onboard compute limits are not yet tested. The policy also lacks formal safety or constraint guarantees. Future directions include hardware validation with varied sensors, rigorous benchmarking of slack–taut transitions, extension to contacts and time-varying cable length, testing on aggressive infeasible references, integration of safety wrappers (e.g., CBF shields) and online $(m_{P},l)$ identification, exploration of meta-RL and few-shot adaptation for out-of-distribution payloads and environments, scaling to multi-UAV cooperative transport, and developing analytical links (e.g., input-to-state stability, robustness margins) between the learned policy and control guarantees.